OpenStudy (chaotic_butterflies):

Graph x^2 - y^2 = 36

OpenStudy (chaotic_butterflies):

@Nnesha

OpenStudy (chaotic_butterflies):

I'm pretty sure this is suppossed to be a hyperbola, but I don't understand how to graph it

OpenStudy (whovianchick):

type your equation into desmos.com and it does it for you!

OpenStudy (chaotic_butterflies):

I don't want an online graph, I want to know how to

OpenStudy (whovianchick):

https://www.desmos.com/calculator

OpenStudy (whovianchick):

ohhhh gotcha

OpenStudy (whovianchick):

It's been a while since I've done that sorry lol

OpenStudy (chaotic_butterflies):

@Vocaloid

OpenStudy (chaotic_butterflies):

when I factor out x^2 - y^2 = 36, I get: (x-y) (x+y) = 6^2

Nnesha (nnesha):

thats not gonna work you need to find vertices foci , and the equation for a line

OpenStudy (chaotic_butterflies):

How do I approach that?

Nnesha (nnesha):

first of all is it horizontal or vertical ? do you know ?

Nnesha (nnesha):

^without looking at the graph ) \[\large\rm \frac{ (x-h)^2 }{ a^2 }-\frac{(y-k)^2}{b^2}=1\] if x comes first then its gonna be horizontal transverse axis

Nnesha (nnesha):

or if the y comes first \[\large\rm \frac{ (y-k)^2 }{ a^2 }-\frac{(x-h)^2}{b^2}=1\]